OpenStudy (anonymous):

i need mode median mean 10,6,8,6,1,4

OpenStudy (saifoo.khan):

mean = add them then divide by number of terms

OpenStudy (anonymous):

but i dont know how to divide

OpenStudy (anonymous):

sum the elements, and divide by the number of elements there are.

OpenStudy (anonymous):

mode is most often i believe so that would be 6. median is you put the numbers in numerical order and it is the one in the middle which would be 6 i think because the order is 1,4,6,6,8,10 and 6 is in the middle. and the mean would be 1+4+6+6+8+10=35 then 35/6= 35/6 as a fraction or 5.8. I think this is correct